The precipitation of aluminum hydroxide, Al(OH), (K 1.3 × 10-33), is sometimes used to purify water. Part A At what pH will precipitation of A1(OH)3 begin if 6.80 lb of aluminum sulfate, Alb (SO4)3. is added to 1550 gallons of water with a negligible change in volume)? Express your answer numerically to two decimal places. Hints pH= 3.88 Submit My Answers Give Up Correct Part B Approximately how many pounds of calcium oxide, CaO, must be added to the water to achieve this pH? Express your answer in pounds using two significant figures. Hints lb Submit My Answers Give Up Provide Feedback Continue
Answer #1

Solution -As you obtained pH as 3.88

poH= 10.12

=>OH-=10^-POH= 10^-10.12 =7.6*10^-11 m

Volume of solution=1550 gallon

=1550*4*0.946

=5865.2L

Moles of OH^-=volume*(OH^-)

7.6*10^-11* 5865.2

=4.45*10^-7 mol

Moles of CaO=1/2*Moles of OH^-

=0.5*4.45*10^-7

=2.2*10^-7 mol

Mass of CaO= moles*molar mass of CaO

=2.2*10^-7* 56.08

=1.2*10^-5g

=1.2*10^-5/453.6

= 2.64*10^-8 pounds